Define elasticity of supply and find the price from the given statement:


If Es of a good is 2 and a firm supplies 200 units at price of Rs 8 per unit, then at what price will the firm supply 250 units.


1
Expert's answer
2021-05-26T10:30:18-0400
"E=\frac {Q_2-Q_1}{p_2-p_1}\times \frac {p_1}{Q_1}"

"2=\frac{250-200}{p_2-8}\times \frac{8}{200}"

"p_2=9"